Fermi National Laboratory


Accelerator Update

Friday July 30
The day shift began with Operations monitoring store 3685 and the Recycler stash, with the Antiproton Source stacking, with the Meson MCenter and MTest experiments taking mean, but with no beam to MiniBooNE due to Horn work.

At 10:21 AM, the Tevatron (TeV) suffered a sector E1 quench. Investigations revealed that the probable cause was due to a quadrupole power supply's (C:B0Q6) output changing. The investigation continued.

MiniBooNE experts reported at 11:13 AM that the Horn was being reconnected. However, they ran with the Horn off.

An MTest experimenter said they were finished with beam at 11:16 AM.

Operations prepared the D-Zero (D0) experiment for access at 12:07 PM.

A Main Injector (MI) expert began a slipstacking study at 1:45 PM.

At 3 PM, Operations disabled the MI beam so technicians could repair a LCW leak on a quadrupole power supply at MI10.

Operations put the TeV through a dry squeeze at 3:23 PM.

Power supply techs at MI10 reported at 3:42 PM that they had temporarily patched a leak on the bus; a full repair would take six hours.

Power supply techs finished their work on B0Q6. Operations put the TeV through a dry squeeze at 4:14 PM to verify the supply was working. The magnet B0Q6 quenched at 4:23 PM. The techs reported at 4:44 PM that they replaced the power supply's regulation card.

The MI expert changed her study to a coalescing study. She completed her study at 5:12 PM.

Operations began sending beam to MiniBooNE around 5:15 PM.

At 7:36 PM, during the TeV's first ramp, the B0Q3 magnet quenched. The cause for the quench was uncertain. The best theory was that the cryo system wasn't quite ready, but there was nothing obvious to prove that. The investigation continued.

Operations put the TeV through a dry squeeze at 9:23 PM and through a wet squeeze at 10:26 PM.

Operations put the TeV into shot setup at 10:52 PM. (Beam to Switchyard (SY) is disabled during shot setup.)

Saturday July 31
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash, and with MiniBooNE and MCenter experiments taking beam.

Operations established store 3692 at 12:34 PM with an initial luminosity of 79.66E30.

Pbar experts flew wires, and then resumed stacking at 12:55 AM.

Operations reestablished beam to MCenter at 1:14 AM.

At 1:33 AM, a Booster RF station (BRF13) tripped off. It reset okay.

At 2:15 AM, BRF18 tripped off. It reset okay.

Operations reported at 6 AM that stacking and beam to SY had been erratic due to Time Line Generator (TLG) and tuning problems.

Operations established beam to MTest at 6:28 AM.

Operations reported at 120:23 AM that a Booster power supply (EAPS) had tripped off twice due to problems with the TLG.

At 12:31 PM, a Linac RF station (LRF2) tripped off due to its RF water pump tripping off. This held off all down stream beam. With the help of an expert, operators had the station back up by 3:09 PM. When LRF2 went down, Linac techs came in and began changing the failing power amplifier (PA) for LRF5.

At 1:26 PM, the TeV suffered a four-house quench starting at sector E1.

Linac techs reported at 4:10 PM that they had completed the LRF5 PA change.

TeV experts reported at 5:21 PM that B0Q3 had tripped off due to a ground fault and was the cause of the quench.

D0 experimenters made a twenty-minute access in to their collision hall at 6:39 PM.

TeV experts reported at 7:15 PM that they hipotted B0Q3 and found it was shorted.

Sunday August 1
The midnight shift began with the TeV off due to a B0Q3 ground fault, with Operations monitoring the stash, with Pbar stacking, and with MiniBooNE, MCenter, and MTest taking beam.

TeV experts and MCR operators accessed CDF and reported at 1:32 AM that they could not visually inspect B0Q6 due to shielding. However, they saw something of interest, a frost mark on one block of shielding (see picture below). The shielding will have to be removed. Experts decided to get together in the morning and determine the next step.

MTest experimenters reported at 12:24 AM that they were finished with beam for the night.

Operations reported at 12:40 AM that they were trying to resolve a Pbar Accumulator emittance problem that had reduced the stacking rate.

At 8:59 AM, LRF4 tripped off due to a small LCW leak on its RF pump. It took a bit of work to make up the lost water. When operators tried to turn on the station at 11:34 AM there were still problems.

At 11:38 AM, LRF2 went down due to a LCW pump problem. Operators had it back on by 12:23 PM.

At 12:28 PM, operators reported that the LRF4 tube was sparking. Operations called in Linac techs.

At 2:20 PM, KRF5 tripped off. It reset okay.

Linac experts and MCR operators reported at 2:24 PM that they had found a LCW leak in the LRF4 filament power supply. This hose leak was a slow dripper that blew. Linac techs had the station back up at 3:18 PM.

Recycler experts began stashing antiprotons from Pbar, without a tune up, at 2:26 PM. Pbar resumed stacking after the transfers ended.

Operations halted stacking at 4:31 PM so that Pbar experts could conduct a reverse proton study.

TeV experts reported at 5:07 PM that they found the B0Q3 fault. After the shielding had been removed they found a leaking Kautzky valve. This had caused an ice ball that reached a broken voltage tap cover for the B0Q3 magnet (See picture below).

Techs repaired the Kautzky valve and after cleaning the voltage tap, the magnet checked out okay.

At 6:10 PM, LRF3 tripped off. It reset okay.

At 6:20 PM, the Linac Klystron Buncher and dipole power supply tripped off. They both reset okay.

The Collider Detector at Fermilab (CDF) reported at 8:43 PM that the TeV B0Q3 shielding had been replaced.

The Cryo Coordinator reported at 8:54 PM that techs had started the B0 cryo system cool-down.

An expert hipotted the B0Q3 power supply at 8:51 PM and it checked out good.

Pbar experts completed their reverse proton study at 11:44 PM.

Pbar resumed stacking at 11:58 PM.

Monday August 2
The midnight shift began with the TeV turn on and checkout beginning, with Operations monitoring the stash, with Pbar stacking, and with beam to MiniBooNE, but with no beam to SY.

At 2:08 AM, the TeV turn on stopped due to a power supply (BQ0) not coming on. Operators called in an expert. The expert reported at 3:04 AM that he had found and replaced a bad relay.

At 3 AM, a TeV RF station (TRF8) tripped off due to a heater ground fault. Operators managed to get the station back on thirty minutes later.

Operations put the TeV through a wet squeeze around 3:30 AM. This study lasted for an hour.

Operations established a 4x0 study store at 7:49 AM.

Plans for Monday
The plan for today is to go into shot setup as soon as possible after the studies and then to continue with stacking and storing.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/2/2004   email Fermilab